# Barbie: Text to Barbie-Style 3D Avatars

### v2

@@ Abstract
Recent advances in text-guided 3D avatar generation have made substantial progress by distilling knowledge from diffusion models. Despite the plausible generated appearance, existing methods cannot achieve fine-grained disentanglement or high-fidelity modeling between inner body and outfit. In this paper, we propose Barbie, a novel framework for generating 3D avatars that can be dressed in diverse and high-quality Barbie-like garments and accessories. Instead of relying on a holistic model, Barbie achieves fine-grained disentanglement on avatars by semantic-aligned separated models for human body and outfits. These disentangled 3D representations are then optimized by different expert models to guarantee the domain-specific fidelity. To balance geometry diversity and reasonableness, we propose a series of losses for template-preserving and human-prior evolving. The final avatar is enhanced by unified texture refinement for superior texture consistency. Extensive experiments demonstrate that Barbie outperforms existing methods in both dressed human and outfit generation, supporting flexible apparel combination and animation. The code will be released for research purposes. Our project page is: https://xiaokunsun.github.io/Barbie.github.io/.
@@ 1 Introduction
In recent years, the creation of 3D digital humans has garnered significant attention due to its widespread applications in AR/VR. These applications yield the generated 3D avatars to have exquisite geometry, lifelike appearance, ultra-high diversity, and a faithful disentanglement of the human body and apparel. Usually, manually sculpting holistic 3D virtual humans is highly labor-intensive and time-consuming. Although progress (Ranjan et al. 2018; Ma et al. 2020; Chen et al. 2022) have been made on learning-based automatic human generation, these methods require large amounts of 3D human data for training, which greatly limits their application scope. Thanks to rapid progress on text-to-image (Rombach et al. 2022) and text-to-3D (Poole et al. 2022), leveraging natural language inputs for avatar generation has become increasingly popular recently.
@@
The aforementioned text-to-avatar works can be roughly divided into two categories: generating (1) a holistic avatar and (2) disentangled models on body and apparel. Works on the former one (Hong et al. 2022; Kolotouros et al. 2024; Liao et al. 2024) are based on text-to-3D generation frameworks (Poole et al. 2022; Chen et al. 2023; Lin et al. 2023), and combined with human prior knowledge (Loper et al. 2015; Pavlakos et al. 2019; Alldieck, Xu, and Sminchisescu 2021) to improve the geometry fidelity of the generated results. Nevertheless, these works treat the human body, clothes, and accessories as a holistic model, which results in a loss of flexibility on applications such as virtual try-on and apparel transfer. As illustrated in Fig. 1-(a), these works cannot properly control garments and accessories with prompt. Consequently, a series of disentangled text-to-avatar works (Huang et al. 2023b; Wang et al. 2023a; Dong et al. 2024) have emerged. These methods employ multi-stage optimization strategies to model the body and clothes separately. However, these methods adopt a single general diffusion model to guide both body and garment generation, resulting in domain-specific fidelity loss on geometry or texture details. Furthermore, as illustrated in Fig. 1-(b), they fail to generate diverse and realistic outfits such as necklaces, glasses, hats, watches, or other accessories. The reason behind is two-fold: these methods lack of (1) fine-grained decomposition on avatar representation; (2) suitable constraints for optimization to generate in-domain components. In addition to this, generating fine-grained disentangled 3D avatars with realistic and text-aligned bodies, as well as garments and accessories, is still an open problem.
Figure 2: Using only texts as input, our method can generate Barbie-style 3D avatars. ”Barbie-style” refers to (1) the generated 3D avatars exhibiting high quality and diversity; (2) the generated body, garments, and accessories being disentangled; and (3) the generated results supporting seamless apparel composition and user-friendly animation, similar to Barbie dolls.
In this paper, we propose Barbie, a novel framework for generating Barbie-style 3D avatars that can be dressed in high-fidelity and diverse outfits, similar to the design of Barbie dolls. As shown in Fig. 2, the avatars generated by Barbie have exquisite geometry and textures, and meanwhile wear various generated high-quality clothes and accessories that can be freely combined. Instead of generating bodies and all kinds of outfits guided by a single text-to-image (T2I) model, Barbie suitably involves different expert diffusion models to guarantee the domain-specific fidelity. Concretely, we first initialize different components of an avatar with semantic-aligned human prior to get disentangled representations. Then, each representation is optimized by corresponding expert T2I models to generate in-domain geometry and texture. During optimization, we propose human-prior evolving loss to enhance the diversity and description alignment for body generation, and template-preserving loss to suppress the artifacts of generated outfits. Finally, to balance the effect of different expert models, we propose an unified texture refinement to jointly optimize the composed avatar for more realistic appearance generation. In this way, the generated 3D avatars are disentangled and interactive with high-fidelity body, garments, and accessories, which is actually ‘Barbie-style’.
@@ 2 Related Work
Text to 3D Content Generation. Motivated by the success of text-guided image generation, many works have focused on generating 3D content from textual prompts in a zero-shot manner. Early methods (Sanghi et al. 2022; Jain et al. 2022; Mohammad Khalid et al. 2022) use the CLIP model (Radford et al. 2021) to constrain the underlying 3D representations to align with the text descriptions. Nevertheless, the quality of created content is not satisfactory due to CLIP’s limited generation capabilities. DreamFusion (Poole et al. 2022) then propose a score distillation sampling (SDS) loss, which enables more high-fidelity 3D content generation by utilizing a more powerful pre-trained T2I model (Rombach et al. 2022). Subsequently, a series of excellent works have emerged to boost the quality of text-to-3D generation through improvements on 3D representations (Chen et al. 2023; Yi et al. 2024), optimization strategies (Lin et al. 2023; Tang et al. 2023), SDS loss (Wang et al. 2024; Liang et al. 2024), and diffusion models (Li et al. 2023c; Qiu et al. 2024; Shi et al. 2023). However, these methods do not fully utilize the rich prior of the human body or generate realistic and high-quality 3D avatars.
@@
Text to Holistic 3D Avatar Generation. AvatarCLIP (Hong et al. 2022) first realizes the zero-shot generation of 3D digital humans from text prompts by integrating human prior knowledge into Neus (Wang et al. 2021). With the introduction of SDS loss (Poole et al. 2022), many subsequent works (Jiang et al. 2023; Cao et al. 2024; Kolotouros et al. 2024; Mendiratta et al. 2023) combine it with parametric human models (Loper et al. 2015; Pavlakos et al. 2019; Alldieck, Xu, and Sminchisescu 2021), significantly improving the fidelity of the generated 3D digital humans. In addition, AvatarVerse (Zhang et al. 2024) and DreamWaltz (Huang et al. 2024b), inspired by ControlNet (Zhang, Rao, and Agrawala 2023), greatly enhance the 3D consistency of the generated results by adopting conditional diffusion models. TADA (Liao et al. 2024) and X-Oscar (Ma et al. 2024) propose using a subdivided SMPL-X model (Pavlakos et al. 2019) instead of implicit NeRF (Mildenhall et al. 2020) and Neus (Wang et al. 2021) to represent 3D digital humans. HumanNorm (Huang et al. 2024a) and SeeAvatar (Xu, Yang, and Yang 2023) achieve impressive geometric details by combining DMTet (Shen et al. 2021) with decoupled optimization strategies. Recently, HumanGaussian (Liu et al. 2024) and GAvatar (Yuan et al. 2024) introduced 3D Gaussian Splatting (Kerbl et al. 2023) to create fast-rendering high-quality 3D avatars. However, these approaches typically treat the human body, garments, and accessories as a holistic model, which loses the flexibility and controllability of generation for further applications, such as the composition of clothing.
@@
Text to Disentangled 3D Avatar Generation. To realize controllable avatar generation, a series of decoupled approaches have been proposed that separately model the human body and clothing through multi-stage optimization. Humancoser (Wang et al. 2023b) and AvatarFusion (Huang et al. 2023b) using two radiation fields (Mildenhall et al. 2020; Wang et al. 2021) to represent the human body and all clothes, respectively. LAGA (Gong et al. 2024) and TELA (Dong et al. 2024) employ a layer-wise framework, representing each piece of clothing as an independent layer, further enhancing the flexibility and controllability of the generation. SO-SMPL (Wang et al. 2023a) replaces implicit representations (Mildenhall et al. 2020; Wang et al. 2021) with a more explicit SMPL-X (Pavlakos et al. 2019) representation. Nonetheless, these works employ a single general T2I model to guide both body and clothing generation, leading to a loss of geometry or texture fidelity in specific domains. Furthermore, they are unable to generate diverse and realistic outfits, such as necklaces, glasses, hats, watches, or other accessories.
@@
In contrast to previous methods, the digital human generated by Barbie not only has fine geometry and realistic appearance but also wears multiple separable high-quality clothes and accessories. We summarize the main differences between our approach and related approaches in Fig. 1.
@@
Figure 3: Barbie consists of three stages: (a) Using human-specific diffusion models and self-evolving human prior loss, Barbie generates a plausible and high-fidelity basic human body; (b) Initializing with the semantic-aligned template mesh, Barbie creates high-quality clothes and accessories piece by piece using object-specific diffusion models and template-preserving loss; (c) Finally, Barbie jointly fine-tunes the assembled avatar to further enhance the texture harmony and consistency.
3 Methodology
@@ 3.1 Preliminaries
SMPL-X (Pavlakos et al. 2019) is a parametric human model that represents the full-body shape, pose, and expression using a small number of parameters. Given a set of shape parameters $\beta$, body pose parameters ${\theta}_{\text{body}}$, jaw pose parameters ${\theta}_{\text{face}}$, finger pose parameters ${\theta}_{\text{hand}}$, and expression parameters $\psi$, it can generate a 3D human mesh ${M}$.
Score Distillation Sampling (Poole et al. 2022) leverages a pre-trained T2I model to guide the 3D representation to align with the input text. Given a text prompt $y$, a 3D representation with parameters $\theta$, and a diffusion model with parameters $\phi$, the SDS loss is defined as follows:
$\nabla_{\theta}\mathcal{L}_{SDS}=\mathbb{E}_{t,\epsilon}\left[w(t)({\epsilon}_{\phi}(x_{t};y,t)-\epsilon)\frac{\partial x}{\partial\theta}\right],$ (1)
@@
DMTet (Shen et al. 2021) is a hybrid representation consisting of an implicit signed distance function (SDF) and a differentiable marching tetrahedra layer. The SDF can be efficiently represented by an MLP with multi-resolution hash encoding (Müller et al. 2022).
@@ 3.2 Overview
Given a prompt, Barbie aims to generate a disentangled 3D digital human dressed in high-fidelity and diverse garments and accessories, similar to Barbie dolls. Fig. 3 shows the overview of the pipeline with three crucial stages. To guarantee the quality of generated body and outfits, we employ targeted expert diffusion models with specific regularization to produce a high-fidelity and reasonable human body (Sec. 3.3) and outfits (Sec. 3.4). The final composed avatar is jointly fine-tuned with unified texture refinement to achieve a more harmonious and consistent appearance (Sec. 3.5).
@@ 3.3 Human Body Generation
Human Body Initialization. Following HumanNorm and RichDreamer, we adopt DMTet as our 3D representation. Since DMTet is highly sensitive to initialization, we use SMPL-X mesh to build an accurate initial input. Concretely, we employ a differentiable renderer (Laine et al. 2020) and SDS loss to optimize the shape parameters $\beta$ according to the input base human body description, thereby determining the basic body shape of the digital human. This initialized mesh $M_{\text{init}}$ will be used in subsequent stages to provide rich human prior knowledge and ensure a semantic-aligned representation for the outfit.
Human Body Geometry Modeling. We then learn to optimize the initialized DMTet into disentangled geometry and texture. As discussed in introduction, using a single general T2I model struggles to provide domain-specific constraints for creating realistic human body or outfits. Therefore, we employ the human-specific diffusion models in HumanNorm (Huang et al. 2024a) which is fine-tuned on high-fidelity 3D human data for detailed body modeling.
@@
Self-Evolving Human Prior Loss. Although Eqs. (2) and (3) leads to detailed human body generation, the overly strong generative prior of human-specific diffusion models may urge the generated results to overfit the input text, resulting in unnatural geometry and exaggerated body proportions. Introducing the constraints from parametric human models is an intuitive solution, but it loses the diversity and fine details. To solve this problem, we draw inspiration from SeeAvatar (Xu, Yang, and Yang 2023) and propose a self-evolving human prior loss. In particular, we periodically fit the initial mesh $M_{\text{init}}$ to the generated DMTet, which is formulated as:
$\mathcal{L}_{\text{prior}}=\sum_{p\in P}\left\|s_{\theta_{\text{h}}}(p)-s_{\hat{M}_{\text{init}}}(p)\right\|_{2}^{2},$ (4)
where $s_{\theta_{\text{h}}}$ and $s_{\hat{M}_{\text{init}}}$ are the SDF functions of the generated body DMTet and the periodically evolved mesh from $M_{\text{init}}$, respectively. $P$ is a set of randomly sampled points in space. $s_{\theta_{\text{h}}}$ is optimized every step, while $s_{\hat{M}_{\text{init}}}$ is optimized every 1000 steps. In this way, $s_{\theta_{\text{h}}}$ is regularized by the reliable human prior from $\hat{M}_{\text{init}}$, while $\hat{M}_{\text{init}}$ is self-evolved to fit the current avatar geometry by learnable vertex displacements that represent geometric details. During the evolution process, $\hat{M}_{\text{init}}$ gradually captures rich geometric features, such as muscle lines, providing reliable but diverse human priors for subsequent geometry creation. Compared with simply using SMPL-X SDF constraints, our self-evolving human prior loss enhances the limited parametric model, achieving a balance between diversity and rationality of geometry generation. In contrast to SeeAvatar, our loss retains the topological structure of SMPL-X, supporting the subsequent outfit initialization, composition, and animation.
@@
$\mathcal{L}_{\text{hum-geo}}=\mathcal{L}^{\text{hn}}_{SDS}+\mathcal{L}^{\text{hd}}_{SDS}+\lambda_{\text{prior}}\mathcal{L}_{\text{prior}}.$ (5)
Human Body Texture Modeling. Given the human mesh generated from the previous stage, we can create realistic and normal-aligned textures using ${\phi}_{\text{hc}}$ with loss as:
$\nabla_{{\theta}_{\text{h}}}\mathcal{L}^{\text{hc}}_{SDS}=\mathbb{E}_{t,\epsilon}\left[({\epsilon}_{{\phi}_{\text{hc}}}(c^{\text{h}}_{t};n^{\text{h}},y_{\text{h}},t)-\epsilon)\frac{\partial c^{\text{h}}}{\partial{\theta}_{\text{h}}}\right],$ (6)
where $c^{\text{h}}$ represents the rendered color image of the generated human body. It’s worth mentioning that experiments using Physically Based Rendering (PBR) lead to unnatural or unrealistic appearances. Empirically, we adopt simple RGB rendering, and the experimental comparison are available in the supplementary material.

### v5

@@ Abstract
Recent advances in text-guided 3D avatar generation have made substantial progress by distilling knowledge from diffusion models. Despite the plausible generated appearance, existing methods cannot achieve fine-grained disentanglement or high-fidelity modeling between inner body and outfit. In this paper, we propose Barbie, a novel framework for generating 3D avatars that can be dressed in diverse and high-quality Barbie-like garments and accessories. Instead of relying on a holistic model, Barbie achieves fine-grained disentanglement on avatars by semantic-aligned separated models for human body and outfits. These disentangled 3D representations are then optimized by specialized expert models to guarantee the domain-specific fidelity. To balance geometry diversity and reasonableness, we propose a series of losses for template-preserving and human-prior evolving. The final avatar is enhanced by unified texture refinement for superior texture consistency. Extensive experiments demonstrate that Barbie outperforms existing methods in both dressed human and outfit generation, supporting flexible apparel combination, editing, and animation.
Figure 1: Using only texts as input, our method can generate Barbie-style 3D avatars. “Barbie-style” refers to: (1) high-quality and diverse generated 3D avatars; (2) disentangled generation of the body, garments, and accessories; and (3) support for seamless apparel composition, editing, and user-friendly animation, similar to Barbie dolls.
@@ 1 Introduction
In recent years, the creation of 3D digital humans has garnered significant attention due to its widespread applications in AR/VR. These applications yield the generated 3D avatars to have exquisite geometry, lifelike appearance, ultra-high diversity, and a faithful disentanglement of the human body and apparel. Usually, manually sculpting holistic 3D virtual humans is highly labor-intensive and time-consuming. Although progress [61, 48, 5] has been made on learning-based automatic human generation, these methods require large amounts of 3D human data for training, which greatly limits their application scope. With recent advancements in text-to-image [65] and text-to-3D [58] models, leveraging natural language inputs for avatar generation has become increasingly popular.
The aforementioned text-to-avatar works can be roughly divided into two categories: generating (1) a holistic avatar and (2) disentangled models on body and apparel. Works on the former one [23, 32, 42] are based on text-to-3D generation frameworks [58, 4, 43], and combined with human prior knowledge [46, 57, 1] to improve the geometry fidelity of the generated results. However, these works treat the human body, clothes, and accessories as a holistic model, which results in a loss of flexibility on applications such as virtual try-on and garment transfer. As illustrated in Fig. 2-(a), these works cannot properly control garments and accessories with prompt. Consequently, a series of disentangled text-to-avatar works [25, 72, 14] have emerged. These methods employ multi-stage optimization strategies to model the body and clothes separately. However, these methods adopt a single general diffusion model to guide both body and garment generation, resulting in domain-specific fidelity loss on geometry or texture details. Furthermore, as illustrated in Fig. 2-(b), they fail to generate diverse and realistic outfits such as necklaces, glasses, hats, watches, or other accessories. The reason behind is two-fold: these methods lack of (1) fine-grained decomposition on avatar representation; (2) suitable constraints for optimization to generate in-domain components. In addition to this, generating fine-grained disentangled 3D avatars with realistic and text-aligned bodies, as well as garments and accessories, is still an open problem.
@@
In this paper, we propose Barbie, a novel framework for generating Barbie-style 3D avatars that can be dressed in high-fidelity and diverse outfits, similar to the design of Barbie dolls. As shown in Fig. 1, the avatars generated by Barbie have exquisite geometry and textures, and meanwhile wear various generated high-quality clothes and accessories that can be freely combined and edited. Instead of generating bodies and all kinds of outfits guided by a single text-to-image (T2I) model, Barbie suitably involves different expert diffusion models to guarantee the domain-specific fidelity. Concretely, we first initialize different components of an avatar with semantic-aligned human prior to get disentangled representations. Then, each representation is optimized by corresponding expert T2I models to generate in-domain geometry and texture. During optimization, we propose SMPLX-evolving prior loss to enhance the diversity and description alignment for body generation, and a template-preserving loss to suppress the artifacts of generated outfits. Finally, to balance the effect of different expert models, we propose an unified texture refinement to jointly optimize the composed avatar for more realistic appearance generation. In this way, the generated 3D avatars are disentangled and interactive with high-fidelity body, garments, and accessories, which is actually “Barbie-style”.
@@ 2 Related Work
Text to 3D Content Generation. Motivated by text-guided image generation, many works have focused on zero-shot 3D content generation from text prompts. Early methods [66, 28, 52] use the CLIP model [60] to align 3D representations with text descriptions, but the quality is unsatisfactory. DreamFusion [58] introduces a score distillation sampling (SDS) loss, enabling higher-fidelity 3D content by leveraging a stronger pre-trained T2I model [65]. Later works improve text-to-3D generation through advancements in 3D representations [4, 77], optimization strategies [43, 70], SDS loss [75, 41], and diffusion models [40, 59, 69]. However, these methods do not fully utilize the rich prior of the human body or generate realistic and high-quality 3D avatars.
@@
Text to Holistic 3D Avatar Generation. AvatarCLIP [23] first realizes the zero-shot generation of 3D digital humans from text prompts by integrating human prior knowledge into Neus [73]. With the introduction of SDS loss [58], many subsequent works [29, 3, 32, 50] combine it with parametric human models [46, 57, 1], significantly improving the fidelity of the generated 3D digital humans. In addition, AvatarVerse [79] and DreamWaltz [27], inspired by ControlNet [80], greatly enhance the 3D consistency of the generated results by adopting conditional diffusion models. TADA [42] and X-Oscar [49] propose using a subdivided SMPL-X model [57] instead of implicit NeRF [51] and Neus [73] to represent 3D digital humans. HumanNorm [26] and SeeAvatar [76] achieve impressive geometric details by combining DMTet [68] with decoupled optimization strategies. Recently, HumanGaussian [44] and GAvatar [78] introduced 3D Gaussian Splatting [31] to create fast-rendering high-quality 3D avatars. However, these approaches typically treat the human body, garments, and accessories as a holistic model, which loses the flexibility and controllability of generation for further applications, such as the composition and editing of clothing.
@@
Text to Disentangled 3D Avatar Generation. To realize controllable avatar generation, a series of decoupled approaches have been proposed that separately model the human body and clothing through multi-stage optimization. Humancoser [74] and AvatarFusion [25] using two radiation fields [51, 73] to represent the human body and all clothes, respectively. LAGA [16] and TELA [14] employ a layer-wise framework, representing each piece of clothing as an independent layer, further enhancing the flexibility and controllability of the generation. SO-SMPL [72] replaces implicit representations [51, 73] with a more explicit SMPL-X [57] representation. However, these works employ a single general T2I model to guide both body and clothing generation, leading to a loss of geometry or texture fidelity in specific domains. Furthermore, they are unable to generate diverse and realistic outfits, such as necklaces, glasses, hats, watches, or other accessories.
@@
In contrast to previous methods, the digital human generated by Barbie not only has fine geometry and realistic appearance, but also wears multiple separable high-quality clothes and accessories. We summarize the main differences between Barbie and related approaches in Fig. 2.
3 The Proposed Method
@@
Figure 3: Barbie consists of three stages: (a) Using human-specific diffusion models and SMPLX-evolving prior loss, Barbie generates a plausible and high-fidelity basic human body; (b) Initializing with the semantic-aligned template mesh, Barbie creates high-quality clothes and accessories piece by piece using object-specific diffusion models and template-preserving loss; (c) Finally, Barbie jointly fine-tunes the assembled avatar to further enhance the texture harmony and consistency.
@@ 3.1 Preliminaries
SMPL-X [57] is a parametric human model that represents the full-body shape, pose, and expression using a small number of parameters. Given a set of shape parameters $\beta$, body pose parameters ${\theta}_{\text{body}}$, jaw pose parameters ${\theta}_{\text{face}}$, finger pose parameters ${\theta}_{\text{hand}}$, and expression parameters $\psi_{\text{exp}}$, it can generate a 3D human mesh ${M}$.
Score Distillation Sampling [58] leverages a pre-trained T2I model to guide the 3D representation to align with the input text. Given a text prompt $y$, a 3D representation with parameters $\theta$, and a diffusion model with parameters $\phi$, the SDS loss is defined as follows:
$\nabla_{\theta}\mathcal{L}_{SDS}=\mathbb{E}_{t,\epsilon}\left[w(t)({\epsilon}_{\phi}(x_{t};y,t)-\epsilon)\frac{\partial x}{\partial\theta}\right],\vspace{-0.2cm}$ (1)
@@
DMTet [68] is a hybrid representation consisting of an implicit signed distance function (SDF) and a differentiable Marching Tetrahedral layer for transfering implicit SDF to a explicit mesh. The SDF can be efficiently represented by an MLP with multi-resolution hash encoding [53].
@@ 3.2 Overview
Given a prompt, Barbie aims to generate a disentangled 3D avatar dressed in high-fidelity and diverse garments and accessories, similar to Barbie dolls. Fig. 3 shows the overview of the pipeline with three crucial stages. To guarantee the quality of generated body and outfits, we employ targeted expert diffusion models with specific regularization to produce a high-fidelity and reasonable human body (Sec. 3.3) and outfits (Sec. 3.4). The final composed avatar is jointly fine-tuned with unified texture refinement to achieve a more harmonious and consistent appearance (Sec. 3.5).
@@ 3.3 Human Body Generation
Human Body Initialization. Following HumanNorm and RichDreamer, we adopt DMTet as our 3D representation. Since DMTet is highly sensitive to initialization, we use SMPL-X mesh to build an accurate initial input. Concretely, we employ a differentiable renderer [33] and SDS loss to optimize the shape parameters $\beta$ according to the input base human body description, thereby determining the basic body shape of the digital human. This initialized SMPL-X mesh $M_{\text{init}}$ will be used in subsequent stages to provide rich human prior knowledge and ensure a semantic-aligned representation for the outfit.
Human Body Geometry Modeling. We then learn to optimize the initialized DMTet into disentangled geometry and texture. As discussed in introduction, using a single general T2I model struggles to provide domain-specific constraints to create realistic human body or outfits. Therefore, we employ human-specific diffusion models in HumanNorm [26] which is fine-tuned with high-fidelity 3D human data for detailed body modeling.
@@
SMPLX-Evolving Prior Loss. Although Eq. 2 and Eq. 3 leads to detailed human body generation, the overly strong generative prior of human-specific diffusion models may urge the generated results to overfit the input text, resulting in unnatural geometry and exaggerated body proportions (see Sec. 4.4). A straightforward approach is to introduce parametric human body models (e.g. SMPL-X) and provide human body priors using the following equation:
$\mathcal{L}_{\text{prior}}=\sum_{p\in P}\left\|s_{\theta_{\text{h}}}(p)-s_{\text{init}}(p)\right\|_{2}^{2},\vspace{-0.3cm}$ (4)
where $s_{\theta_{\text{h}}}$ and $s_{\text{init}}$ represent the SDF functions of the generated body DMTet and the initialized SMPL-X mesh $M_{\text{init}}$, respectively, and $P$ is a set of randomly sampled points in space. However, due to the overly smooth geometry of SMPL-X, this approach may lead to a reduction in diversity and rich detail. Another potential solution is the evolving constraint proposed by SeeAvatar [76] to periodically improve the initial DMTet template. Although producing reasonable geometry, this approach discards the original topology and semantics of SMPL-X, and thus becomes unfriendly to garment initialization or avatar animation.
To address this problem, we propose an SMPLX-evolving prior loss that evolves the SMPL-X geometry by learning its displacement. Specifically, we freeze the SMPL-X model but improve $M_{\text{init}}$ to evolvable $\hat{M}_{\text{init}}$ by adding learnable vertex-wise offsets and periodically fit $\hat{M}_{\text{init}}$ to the current generated body DMTet at every $\delta$ iteration, allowing the learned offsets to model complex body features. The fitting objective is as follows:
$\mathcal{L}_{\text{fit}}=\lambda_{\text{chamf}}\mathcal{L}_{\text{chamf}}+\lambda_{\text{edge}}\mathcal{L}_{\text{edge}}+\lambda_{\text{nor}}\mathcal{L}_{\text{nor}}+\lambda_{\text{lap}}\mathcal{L}_{\text{lap}},$ (5)
where $\mathcal{L}_{\text{chamf}}$ is the Chamfer distance between $\hat{M}_{\text{init}}$ and the current body geometry, $\mathcal{L}_{\text{edge}}$, $\mathcal{L}_{\text{nor}}$ and $\mathcal{L}_{\text{nor}}$ are the edge length regularization, normal consistency loss, and Laplacian smoothness, respectively. Through this SMPLX-evolving process, $\hat{M}_{\text{init}}$ gradually captures rich geometric features (e.g. muscle contours), providing reliable yet diverse priors for subsequent geometry generation. In addition, unlike SeeAvatar, it maintains the semantics and topology of the SMPL-X model, supporting the subsequent outfit initialization, composition, editing, and animation.
@@
$\mathcal{L}_{\text{hum-geo}}=\mathcal{L}^{\text{hn}}_{SDS}+\mathcal{L}^{\text{hd}}_{SDS}+\lambda_{\text{prior}}\mathcal{L}_{\text{prior}}.\vspace{-0.1cm}$ (6)
Human Body Texture Modeling. Given the human mesh generated from the previous stage, we fix it and utilize a texture field ${\psi}_{\text{h}}$ which maps a query position to its color to generate a normal-aligned body appearance. Specifically, we employ a MLP network with multi-resolution hash encoding [53] to construct this field and optimize it using ${\phi}_{\text{hc}}$ with loss as:
$\nabla_{{\psi}_{\text{h}}}\mathcal{L}^{\text{hc}}_{SDS}=\mathbb{E}_{t,\epsilon}\left[({\epsilon}_{{\phi}_{\text{hc}}}(c^{\text{h}}_{t};n^{\text{h}},y_{\text{h}},t)-\epsilon)\frac{\partial c^{\text{h}}}{\partial{\psi}_{\text{h}}}\right],\vspace{-0.1cm}$ (7)
where $c^{\text{h}}$ represents the rendered color image of the generated human body. It’s worth mentioning that experiments using Physically Based Rendering (PBR) lead to unnatural or unrealistic appearances. Empirically, we adopt simple RGB rendering, and the experimental comparisons are available in the Supp. Mat.